OpenStudy (anonymous):

Three multiplied by the sum of 4 and a number is the same as 18 more than the number. Find the number. If n is "the number," which equation could be used to solve for the number? 3n+ 4 =n+ 18 4(3 +n) =n+ 18 3(4 +n) =n+ 18

OpenStudy (radar):

3(4+n)=n+18 do you see why?

OpenStudy (anonymous):

Because you multiply the 3 and 4 then subtract to get just n?

OpenStudy (phi):

Three multiplied by the sum of 4 and a number 3 * ( 4 + n)

OpenStudy (radar):

That sounds like the way to solve for n. 12 +3n = n +18 3n -n = 18-12 2n = 6 n =3 3 would be the number, but do you understand how to translate the given written form to get 3(4+n) = n + 18